• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z.-E. Meziani

    Log entry time 00:05:04 on July23,2003

    Entry number 106293

    keyword=Shift Summary

    Start of swing shift
    16:00 Waiting for beam
    16:35 Moved the target to pickup position for an NMR measurement
    16:50 NMR measurement completed P=37.3%
    Decided not to perform EPR until polarization is higher.
    18:10 Everyone is out of the Hall
    19:30 Start production- run # 2702/22702/ 19:29 - 19:43 2M
    (run with large charge asymmetry, Happex DAQ has a problem)
    19:45 Restarted CODA on right arm.
    19:45 MCC called they lost an RF zone
    20:40 MCC is ready to send beam back.
    Moved the target to empty position to make sure the beam is OK.
    Beam position is ok and Polarized target in position
    21:05 Problem with Happex DAQ still there, ROC24 not connected.
    paged Bryan Moffit/ looking into the problem
    21:08 -21:21 run# 2704 is repeat of run #2702 3M events (good run)
    with beam charge asymmetry of less than 200ppm
    Beam half wave plate in
    MCC need to replace a board for the FSD system.
    22:02 MCC is sending beam back
    22:02-22:17 run # 2705 3M (Happex daq being under investigation)
    22:18 Started to rotate the target spin to 270deg
    22:40 ready for beam. Run 2706 very short because the beam was gone after a
    a short time.
    Beam back at 23:08
    23:08-23:22 run # 2707 (Happex daq is working fine now) 3M
    23:23-23:36 run # 2708 3M
    23:37-23:43 run # 2709 0.16M stopped due to lack of beam.
    23:43 MCC called. Lost RF zone it will take at least half an hour
    ...
    24:00 Shift end (still waiting for beam).